### v2

@@ 2.1 Model Setup
Let $x_{i,t}$ be the $i$-th unit of the cross-sectional panel ${\mathbf{x}}_{t}=(x_{1,t},...,x_{N,t})^{\prime}\in R^{N}$ at time $t$, for example, $x_{i,t}$ can be the stock return of the $i$-th asset at time $t$, we consider the following approximate factor model:
@@
where $x_{i,t}$ is the only observed datum for the $i$-th cross-section at time $t$ ($i=1,...,N$; $t=1,...,T$), ${\mathbf{f}}_{t}$ is an $r$-dimensional vector of common or systematic risk factors, $\boldsymbol{\lambda}_{i}$ is an $r$-dimensional vector of factor loadings, and $e_{i,t}$ is the idiosyncratic component of $x_{i,t}$ that cannot be explained by the common risk factors. We may combine all the cross-sectional units together and write the above equation as
@@
In the next subsection, we briefly introduce the identification issues in Model (3) and the estimators obtained by the method of PCA.
2.2 Identification and PCA Estimation
As the panel of data ${\mathbf{X}}$ is the only observable component in Model (3), and ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$ are latent and unobserved ones, the pair $({\mathbf{F}},\boldsymbol{\Lambda})$ itself is not identifiable since they can be replaced by $({\mathbf{F}}{\mathbf{H}}^{\prime},\boldsymbol{\Lambda}{\mathbf{H}}^{-1})$ for any $r\times r$ non-singular matrix ${\mathbf{H}}$, without altering the equation in (3). Therefore, it is conventional to impose some constraints on the factors and their associated loadings. There are two equivalent sets of identification conditions in the PCA framework. The first one assumes
@@ 2.2 Identification and PCA Estimation
$\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}={\mathbf{I}}_{r}\,\,\text{and}\,\,\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}\,\,\text{is diagonal},$ (4)
and the second one postulates that
$\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}={\mathbf{I}}_{r}\,\,\text{and}\,\,\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}\,\,\text{is diagonal}.$ (5)
See, for example, Bai and Ng (2002) and Fan et al. (2013). Note that ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$ are still not identifiable when the diagonals in (4) or (5) have multiplicity or there exists a diagonal matrix ${\mathbf{H}}$ with the diagonals being $1$ or $-1$ such that $({\mathbf{F}}{\mathbf{H}},\boldsymbol{\Lambda}{\mathbf{H}})$ is still a pair of solutions. In this paper, we do not consider the special case of multiplicities, and the change of signs does not affect our analysis because it gives us an advantage to choose an ${\mathbf{F}}$ with any particular signs to facilitate the interpretations of the estimates. On the other hand, according to the discussion in Bai and Ng (2013), the estimates can be uniquely defined asymptotically under conditions (4) or (5) if the top eigenvalues of the covariance of the panel are distinct. They showed that the rotation matrix ${\mathbf{H}}$ is asymptotically equal to an identity with certain convergence rates. The conditions in (4) and (5) are reasonable when applying the PCA method to estimate factors and loadings. Specifically, the estimated PCA factors $\widehat{\mathbf{F}}$ and loadings $\widehat{\boldsymbol{\Lambda}}$ asymptotically satisfy either (4) or (5), depending on whether PCA is performed on ${\mathbf{X}}{\mathbf{X}}^{\prime}$ (Asymptotic PCA) or ${\mathbf{X}}^{\prime}{\mathbf{X}}$ (PCA method). For further details, see Connor and Korajczyk (1986, 1988) and Chapter 9 of Tsay (2010).
Note that the two estimation procedures are equivalent in the sense that the estimated common component $\widehat{\mathbf{F}}\widehat{\boldsymbol{\Lambda}}^{\prime}$ stays the same, which is due to the nature of the PCA method. In either case, the PCA factors and loadings are difficult to interpret. This poses a challenge for modeling and understanding the dynamic structure and the co-movement in the data, especially for those in finance and economics where understanding the hidden economic mechanism in the data is a main objective to study.
In this article, we adopt the conditions in (4) and pick a specific choice of the rotation matrix based on the APCA approach, ensuring the unique identification of ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$. For example, suppose the true underlying factor and loading matrices are ${\mathbf{F}}^{*}$ and $\boldsymbol{\Lambda}^{*}$, respectively, such that ${\mathbf{X}}={\mathbf{F}}^{*}\boldsymbol{\Lambda}^{*}{{}^{\prime}}+{\mathbf{e}}$, we can show that there exists a rotation matrix ${\mathbf{H}}$ such that the estimated APCA factors $\widehat{\mathbf{F}}$ with $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T={\mathbf{I}}_{r}$ is consistent to ${\mathbf{F}}^{*}{\mathbf{H}}^{\prime}$ (up to a scale). Hence, by redefining ${\mathbf{F}}={\mathbf{F}}^{*}{\mathbf{H}}^{\prime}$ and $\boldsymbol{\Lambda}=\boldsymbol{\Lambda}^{*}{\mathbf{H}}^{-1}$, the model transforms into ${\mathbf{X}}={\mathbf{F}}\boldsymbol{\Lambda}^{\prime}+{\mathbf{e}}$, enabling us to uniquely estimate ${\mathbf{F}}$ and $\boldsymbol{\Lambda}$ through the APCA approach. With this redefined framework, our sparsity assumption will be directly imposed on ${\mathbf{F}}$ in subsequent analysis, as there are no identification issues as discussed above. Our work aims to extend the APCA of Connor and Korajczyk (1986, 1988) to high dimensions using advanced machine learning techniques.
2.3 Sparse Factor Processes
In view of the discussion in Section 2.2, the traditional APCA estimators are difficult to interpret, especially when the dimension is large. In the past decades, various sparse PCA or sparse factor modeling techniques have been proposed in the statistical and econometrics literature. See, for example, Zou et al. (2006), Witten et al. (2009), and Kristensen (2017), among other. Sparse PCA restricts the cardinality of the weight vectors for the PCs so that the PCs are sparse linear combinations of the underlying cross-sectional variables. By setting the factor weights to zero, sparse PCA facilitates the interpretation of the PCs, which improves interpretability without sacrificing the explanatory ability of the PCs or factors. To the best of our knowledge, all the aforementioned approaches assume that the loadings are sparse in the population model, and formulate the estimation of sparse loadings as a regularized optimization problem to estimate principle components with a penalty term, which allows for developing asymptotic inferential theory. Nevertheless, the assumption of sparse population loadings may not be satisfied in many datasets. For example, in finance, the exposure of asset returns to a market factor is universal and non-sparse. In addition, according to Camacho et al. (2020), the sparse PCA method presents disparate and limited performance when modeling sparse data, even in the optimistic case of noise-free data.
@@ 2.3 Sparse Factor Processes
2.4 Sparse APCA: Formulation and Estimation
We first introduce some further notation used in the estimation procedure. For an $N\times 1$ vector ${\mathbf{v}}=(v_{1},...,v_{N})^{\prime}$, $\|{\mathbf{v}}\|_{2}=\sqrt{\sum_{i=1}^{2}v_{i}^{2}}$ is the Euclidean norm and $\|{\mathbf{v}}\|_{0}=\text{card}\{\text{support}({\mathbf{v}})\}$ is the cardinality or number of non-zero elements in ${\mathbf{v}}$. Let $\mathbb{V}$ be the set of $N\times r$ semi-orthogonal matrices and $\mathbb{V}^{\perp}$ be the set consisting of the orthogonal complements of the ones in $\mathbb{V}$. For a matrix ${\mathbf{V}}=({\mathbf{v}}_{1},...,{\mathbf{v}}_{r})\in\mathbb{V}$ and a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$, $\|{\mathbf{V}}\|_{0}\leq{\mathbf{s}}$ implies that $\|{\mathbf{v}}_{i}\|_{0}\leq s_{i}$ for $1\leq i\leq r$.
@@ 2.4.1 One-Factor Case
and consequently, the estimator for $\underline{{\mathbf{f}}}_{1}$ is denoted as $\underline{\widehat{\mathbf{f}}}_{1}=\sqrt{T}\widehat{\mathbf{v}}$ such that $\underline{\widehat{\mathbf{f}}}_{1}^{\prime}\underline{\widehat{\mathbf{f}}}_{1}/T=1$ which satisfies the first condition in (4). Note that the solutions in (6) are the same as those obtained by PCA in Bai and Ng (2002) and Gao and Tsay (2023), among others, if the constraint $\|{\mathbf{v}}\|_{0}\leq s_{1}$ is removed. Furthermore, the problem in (6) is also equivalent to finding the largest eigenvalue associated with an $s_{1}$-sparse eigenvector of the matrix ${\mathbf{S}}$:
@@
In fact, the optimization in (6) is still a sparse PCA problem symbolically, and numerous methods have been developed to obtain sparse eigenvectors during the past decades. See the regularization method with elastic net in Zou et al. (2006) and the penalized matrix decomposition (PMD) algorithm using $\ell_{1}$-penalty in Witten et al. (2009), among others.
@@
where ${\mathbf{S}}$ is the matrix obtained by the noisy observation ${\mathbf{x}}_{t}$, $\boldsymbol{\Sigma}$ is a symmetric matrix whose eigenvectors are the true ones, and ${\mathbf{E}}$ is a random perturbation. The decomposition of (8) is formulated in (S.1) of the online Appendix. If the true largest eigenvector ${\mathbf{v}}_{1}$ of $\boldsymbol{\Sigma}$ is sparse, then it is natural to recover ${\mathbf{v}}_{1}$ from the noisy matrix ${\mathbf{S}}$. This recovery is guaranteed when the error ${\mathbf{E}}$ is of a smaller order using the well-known $\sin\theta$ theorem in Davis and Kahan (1970) under the approximate factor model. Second, for any given vector ${\mathbf{u}}_{0}\in R^{N}$, the power method estimates the first eigenvector by
@@
Under the assumption that $|\widehat{\lambda}|>|\widehat{\lambda}_{i}|$ for $i\geq 2$, it follows that the direction of ${\mathbf{u}}_{k}$ converges to that of $\widehat{\mathbf{v}}_{1}$.
@@
The procedure, presented in Algorithm 1, generates a sequence of intermediate $s_{1}$-sparse eigenvectors $\{{\mathbf{u}}_{1},{\mathbf{u}}_{2},...\}$ from an initial sparse approximation ${\mathbf{u}}_{0}$. At each time stamp $t$, the intermediate vector ${\mathbf{u}}_{t-1}$ is multiplied by ${\mathbf{S}}$, and then the entries are truncated to zeros except for the largest $s_{1}$ entries. The resulting vector is then normalized to unit length. Finally, the estimated first factor process is $\underline{\widehat{\mathbf{f}}}_{1}=\sqrt{T}\widehat{\mathbf{v}}_{1}$, where $\widehat{\mathbf{v}}_{1}$ is the output one in Algorithm 1. Note that the sparsity parameter $s_{1}$ is unknown in practice, and we will propose a cross-validation procedure below to choose such a parameter.
Algorithm 1 can consistently estimate sparse eigenvectors under a factor structure when the corresponding eigenvalues diverge. However, this property may not extend to general covariance matrices. For instance, it is possible to construct a diagonal covariance matrix with sparse eigenvectors, where the algorithm may fail to correctly identify the sparse eigenvectors associated with top eigenvalues for some initial vectors. For instance, consider ${\mathbf{S}}=\mbox{diag}(2,1)$. In this case, the leading eigenvector is ${\mathbf{v}}_{1}=(1,0)^{\prime}$, but the algorithm will incorrectly output ${\mathbf{v}}_{1}=(0,1)^{\prime}$ if the initial vector is ${\mathbf{u}}_{0}=(1,3)^{\prime}$. This occurs because the top eigenvalue is not sufficiently large. However, if ${\mathbf{S}}=\mbox{diag}(4,1)$, where the top eigenvalue is significantly larger—as is typically the case in a factor model—the algorithm correctly identifies the leading eigenvector.
@@ 2.4.2 Multi-Factor Case
In this section, we consider the general case when there are multiple factor processes in Model (3), i.e. $r>1$. In the presence of more than one common factor process, we need to apply Algorithm 1 multiple times and extract all the common factors in a sequential way. Specifically, when $\widehat{\mathbf{v}}_{1}$ is given via Algorithm 1, we subtract the projection on first factor component of the panel and the residual would be $\widetilde{\mathbf{X}}={\mathbf{X}}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{X}}=({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{X}}$. The resulting scaled covariance $\widetilde{\mathbf{S}}_{1}=({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime})$. Then the second eigenvector $\widehat{\mathbf{v}}_{2}$ can be obtained by solving the following optimization problem (11) below:
@@
$\widehat{\mathbf{v}}_{i}=\arg\min_{{\mathbf{v}}^{\prime}{\mathbf{B}}_{i}{\mathbf{v}}=1,\|{\mathbf{v}}\|_{0}\leq s_{i}}{\mathbf{v}}^{\prime}{\mathbf{S}}{\mathbf{v}},$
@@
7:     Update ${\mathbf{B}}_{i+1}={\mathbf{B}}_{i}({\mathbf{I}}_{T}-{\mathbf{q}}_{i}{\mathbf{q}}_{i}^{\prime})$.
8:     Return $\widehat{\mathbf{v}}_{i}\leftarrow\widehat{\mathbf{v}}_{i}/\|\widehat{\mathbf{v}}_{i}\|_{2}$.
9:     $i\leftarrow i+1$.
@@
11:Output: $\{\widehat{\mathbf{v}}_{1},...,\widehat{\mathbf{v}}_{r}\}$;
From the optimization problem in (11) and the procedure in Algorithm 2, we cannot enforce the orthogonality and sparsity at the same time, which is similar to the case in the sparse PCA framework. When there is no $\ell_{0}$-constraint in (6) and (11), we can easily obtain that $\widehat{\mathbf{v}}_{1}$ and $\widehat{\mathbf{v}}_{2}$ are the two normalized and orthogonal eigenvectors associated with the top two eigenvalues of ${\mathbf{S}}$ and, the approach reduces to the traditional APCA method. One of the differences between the one-factor case in (6) and the multi-factor case in (11) is that the estimated factors and loadings in the one-factor case satisfy the conditions in (4), whereas those obtained in the multi-factor case may not. It is important to note that this discrepancy is only a finite sample outcome. Asymptotically, we can demonstrate that the multi-factor case still satisfies the identification conditions in (4). Similar outcomes are also observed in the sparse PCA literature, including Zou et al. (2006) and Witten et al. (2009), as well as in the sparse factor modeling frameworks discussed in Kristensen (2017).
For each sparse vector $\widehat{\mathbf{v}}_{i}$ obtained in Algorithm 2, the estimated factor process is obtained as $\underline{\widehat{\mathbf{f}}}_{i}=\sqrt{T}\widehat{\mathbf{v}}_{i}$. It is not hard to see that the number of iterations is $r$, and the $i$th iteration outputs a sparse estimator (multiplied by $\sqrt{T}$) of the $i$th column of ${\mathbf{F}}$. For the optimization problem in Step 4 of Algorithm 2, by the argument in Lemma A.1 of the Appendix, the matrix ${\mathbf{B}}_{i}$ in each step is a symmetric one. Therefore, we modify the truncated power method in Algorithm 1 and propose the following algorithm to obtain a sparse eigenvector for Step 4 of Algorithm 2.
@@
1:For each ${\mathbf{B}}={\mathbf{B}}_{i}$, perform a singular-value-decomposition (SVD) ${\mathbf{B}}={\mathbf{U}}{\mathbf{D}}{\mathbf{U}}^{\prime}$ and hence ${\mathbf{B}}^{1/2}={\mathbf{U}}{\mathbf{D}}^{1/2}{\mathbf{U}}^{\prime}$. Let ${\mathbf{A}}={\mathbf{B}}^{-1/2}{\mathbf{S}}{\mathbf{B}}^{-1/2}$.
@@
4:     Compute $\widetilde{\mathbf{x}}_{j}=\frac{{\mathbf{B}}^{-1/2}{\mathbf{A}}{\mathbf{x}}_{j-1}}{\|{\mathbf{A}}{\mathbf{x}}_{j-1}\|_{2}}$
5:     Let $l_{j}=\text{supp}(\widetilde{\mathbf{x}}_{j},s)$ be the indices of $\widehat{\mathbf{x}}_{t}$ with the largest $s$ absolute values.
6:     Compute ${\mathbf{x}}_{j}^{*}=\text{Truncate}(\widetilde{\mathbf{x}}_{j},l_{j})$
7:     Let ${\mathbf{x}}_{j}=\frac{{\mathbf{B}}^{1/2}{\mathbf{x}}_{j}^{*}}{\|{\mathbf{B}}^{1/2}{\mathbf{x}}_{j}^{*}\|_{2}}$.
8:     $j\leftarrow j+1$.
9:until ${\mathbf{x}}_{j}$ is convergent

### v3

@@ 2.1 Model Setup
Let $x_{i,t}$ be the $i$-th unit of the cross-sectional panel of time series ${\mathbf{x}}_{t}=(x_{1,t},...,x_{N,t})^{\prime}\in R^{N}$ at time $t$, for example, $x_{i,t}$ can be the stock return of the $i$-th asset at time $t$, we consider the following approximate factor model:
@@
where $x_{i,t}$ is the only observed datum for the $i$-th cross-section at time $t$ (i=1,,Ni=1,...,N; t=1,,Tt=1,...,T), ${\mathbf{f}}_{t}$ is an $r$-dimensional vector of common or systematic risk factors, $\boldsymbol{\lambda}_{i}$ is an $r$-dimensional vector of factor loadings, and $e_{i,t}$ is the idiosyncratic component of $x_{i,t}$ that cannot be explained by the common risk factors. We may combine all the cross-sectional units together and write the above equation as
@@
In Model (3), only the panel $\mathbf{X}$} is observed, while the factor $\mathbf{F}$} and loading $\boldsymbol{\Lambda}$ are latent and unidentifiable up to a rotation: for any invertible $r\times r$ matrix $\mathbf{H}$}, the pair $({\mathbf{F}}{\mathbf{H}}^{\prime},\boldsymbol{\Lambda}{\mathbf{H}}^{-1})$ yields the same model. To ensure identification, two commonly used sets of conditions are adopted in the PCA framework:
@@
$\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}={\mathbf{I}}_{r},\quad\text{and}\quad\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}\text{ is diagonal},\quad\text{or}\quad\frac{1}{N}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}={\mathbf{I}}_{r},\quad\text{and}\quad\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}\text{ is diagonal}.$ (4)
See Bai and Ng (2002), Fan et al. (2013), and Jiang et al. (2023), among others. These identification restrictions help resolve the rotational indeterminacy when the leading eigenvalues of the covariance matrix are distinct; see Bai and Ng (2013). Under either set of conditions, PCA provides consistent estimates of the common component $\widehat{\mathbf{F}}\widehat{\boldsymbol{\Lambda}}^{\prime}$.
In this paper, we adopt the first set of conditions in (4) and apply the Asymptotic Principal Component Analysis (APCA) approach to uniquely identify $\mathbf{F}$} and $\boldsymbol{\Lambda}$. Suppose the true model is ${\mathbf{X}}={\mathbf{F}}^{*}\boldsymbol{\Lambda}^{*^{\prime}}+{\mathbf{e}}$, Jiang et al. (2023) shows that there exists a rotation matrix $\mathbf{H}$} such that the estimated APCA factors $\widehat{\mathbf{F}}$ satisfy $\widehat{\mathbf{F}}^{\prime}\widehat{\mathbf{F}}/T={\mathbf{I}}_{r}$ and are consistent for ${\mathbf{F}}^{*}{\mathbf{H}}^{\prime}$. Redefining ${\mathbf{F}}={\mathbf{F}}^{*}{\mathbf{H}}^{\prime}$ and $\boldsymbol{\Lambda}=\boldsymbol{\Lambda}^{*}{\mathbf{H}}^{-1}$ transforms the model to ${\mathbf{X}}={\mathbf{F}}\boldsymbol{\Lambda}^{\prime}+{\mathbf{e}}$, which allows us to impose sparsity assumptions directly on $\mathbf{F}$} without encountering identification issues.
Our first goal is to extend the APCA framework of Connor and Korajczyk (1986, 1988) to high-dimensional settings using modern machine learning techniques. Moreover, since factors and loadings in approximate factor models are typically difficult to interpret, our second goal, building on the framework of Bai and Ng (2013), is to impose additional sparse structure on the factors. This enables interpretation by linking the estimated factors to specific time events in the data.
2.2 Sparse APCA: Formulation and Estimation
@@
We first introduce some further notation used in the estimation procedure. For an $N\times 1$ vector ${\mathbf{v}}=(v_{1},...,v_{N})^{\prime}$, $\|{\mathbf{v}}\|_{2}=\sqrt{\sum_{i=1}^{2}v_{i}^{2}}$ is the Euclidean norm and $\|{\mathbf{v}}\|_{0}=\text{card}\{\text{support}({\mathbf{v}})\}$ is the cardinality or number of non-zero elements in $\mathbf{v}$}. Let $\mathbb{V}$ be the set of $N\times r$ semi-orthogonal matrices and $\mathbb{V}^{\perp}$ be the set consisting of the orthogonal complements of the ones in $\mathbb{V}$. For a matrix ${\mathbf{V}}=({\mathbf{v}}_{1},\ldots,{\mathbf{v}}_{r})\in\mathbb{V}$ and a vector ${\mathbf{s}}=(s_{1},\ldots,s_{r})^{\prime}$, $\|{\mathbf{V}}\|_{0}\leq{\mathbf{s}}$ means that each column ${\mathbf{v}}_{i}$ satisfies $\|{\mathbf{v}}_{i}\|_{0}\leq s_{i}$ for $1\leq i\leq r$.
@@ 2.2.1 One-Factor Case
and consequently, the estimator for $\underline{{\mathbf{f}}}_{1}$ is denoted as $\underline{\widehat{\mathbf{f}}}_{1}=\sqrt{T}\widehat{\mathbf{v}}$ such that $\underline{\widehat{\mathbf{f}}}_{1}^{\prime}\underline{\widehat{\mathbf{f}}}_{1}/T=1$, which satisfies the first condition in (4). Note that the solutions in (5) are the same as those obtained by PCA in Bai and Ng (2002) and Gao and Tsay (2023), among others, if the constraint $\|{\mathbf{v}}\|_{0}\leq s_{1}$ is removed. Furthermore, the problem in (5) is also equivalent to finding the largest eigenvalue associated with an $s_{1}$-sparse eigenvector of the matrix $\mathbf{S}$}:
@@
The optimization in (5) is still a sparse PCA problem symbolically, and numerous methods have been developed to obtain sparse eigenvectors during the past decades. See the regularization method with elastic net in Zou et al. (2006) and the penalized matrix decomposition (PMD) algorithm using $\ell_{1}$-penalty in Witten et al. (2009), among others.
@@
where $\mathbf{S}$} is the matrix obtained by the noisy observation ${\mathbf{x}}_{t}$, $\boldsymbol{\Sigma}$ is a symmetric matrix whose eigenvectors are the true ones, and $\mathbf{E}$} is a random perturbation. The decomposition of (7) is formulated in (S.1) of the online Supplement. If the true largest eigenvector ${\mathbf{v}}_{1}$ of $\boldsymbol{\Sigma}$ is sparse, then it is natural to recover ${\mathbf{v}}_{1}$ from the noisy matrix $\mathbf{S}$}. This recovery is guaranteed when the error $\mathbf{E}$} is of a smaller order using the well-known $\sin\theta$ theorem in Davis and Kahan (1970) under the approximate factor model. Second, for any given vector ${\mathbf{u}}_{0}\in R^{N}$, the power method estimates the first eigenvector by
@@
Assuming that $|\widehat{\lambda}_{1}|>|\widehat{\lambda}_{i}|$ for $i\geq 2$, it follows that the direction of ${\mathbf{u}}_{k}$ converges to that of $\widehat{\mathbf{v}}_{1}$.
@@
The procedure, presented in Algorithm 1, generates a sequence of intermediate $s_{1}$-sparse eigenvectors $\{{\mathbf{u}}_{1},{\mathbf{u}}_{2},$...\} from an initial sparse approximation ${\mathbf{u}}_{0}$. At each time stamp $t$, the intermediate vector ${\mathbf{u}}_{t-1}$ is multiplied by $\mathbf{S}$}, and then the entries are truncated to zeros except for the largest $s_{1}$ entries. The resulting vector is then normalized to unit length. Finally, the estimated first factor process is $\underline{\widehat{\mathbf{f}}}_{1}=\sqrt{T}\widehat{\mathbf{v}}_{1}$, where $\widehat{\mathbf{v}}_{1}$ is the output one in Algorithm 1. Instead of pre-specifying threshold values as in Johnstone and Lu (2009) and Ma (2013), we rely on the sparsity parameter $s_{1}$, which is unknown in practice. To address this, we will propose a cross-validation procedure to determine the optimal value of $s_{1}$ later.
Algorithm 1 can consistently estimate sparse eigenvectors under a factor structure when the corresponding eigenvalues diverge. However, this property may not extend to general covariance matrices. For instance, it is possible to construct a diagonal covariance matrix with sparse eigenvectors, where the algorithm may fail to correctly identify the leading sparse eigenvectors for some initial vectors. For instance, consider ${\mathbf{S}}=\mbox{diag}(2,1)$. In this case, the leading eigenvector is ${\mathbf{v}}_{1}=(1,0)^{\prime}$, but the algorithm will incorrectly output ${\mathbf{v}}_{1}=(0,1)^{\prime}$ if the initial vector is ${\mathbf{u}}_{0}=(1,3)^{\prime}$. This occurs because the top eigenvalue is not sufficiently large. However, if ${\mathbf{S}}=\mbox{diag}(4,1)$, where the top eigenvalue is significantly larger—as is common in a factor model, the algorithm correctly identifies the leading eigenvector.
@@ 2.2.2 Multi-Factor Case
In this section, we consider the general case where multiple factor processes are present in Model (3), i.e., when $r>1$. We extend the truncated power method to accommodate the multi-factor setting. In the presence of more than one common factor process, we need to apply Algorithm 1 multiple times and extract all the common factors in a sequential way. Specifically, when $\widehat{\mathbf{v}}_{1}$ is given via Algorithm 1, we subtract the projection on first factor component of the panel and the residual would be $\widetilde{\mathbf{X}}={\mathbf{X}}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{X}}=({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{X}}$. The resulting scaled covariance $\widetilde{\mathbf{S}}_{1}=({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime})$. Then the second eigenvector $\widehat{\mathbf{v}}_{2}$ can be obtained by solving the following optimization problem (10) below:
@@
$\widehat{\mathbf{v}}_{i}=\arg\min_{{\mathbf{v}}^{\prime}{\mathbf{B}}_{i}{\mathbf{v}}=1,\|{\mathbf{v}}\|_{0}\leq s_{i}}{\mathbf{v}}^{\prime}{\mathbf{S}}{\mathbf{v}};$
@@
7:  Update ${\mathbf{B}}_{i+1}={\mathbf{B}}_{i}({\mathbf{I}}_{T}-{\mathbf{q}}_{i}{\mathbf{q}}_{i}^{\prime})$;
8:  Return $\widehat{\mathbf{v}}_{i}\leftarrow\widehat{\mathbf{v}}_{i}/\|\widehat{\mathbf{v}}_{i}\|_{2}$;
9:  $i\leftarrow i+1$;
@@
11:Output: $\{\widehat{\mathbf{v}}_{1},...,\widehat{\mathbf{v}}_{r}\}$.
From the optimization problem in (10) and the procedure in Algorithm 2, we cannot enforce the orthogonality and sparsity at the same time, which is similar to the case in the sparse PCA framework. When there is no $\ell_{0}$-constraint in (5) and (10), we can easily obtain that $\widehat{\mathbf{v}}_{1}$ and $\widehat{\mathbf{v}}_{2}$ are the two normalized and orthogonal eigenvectors associated with the top two eigenvalues of $\mathbf{S}$}, and the approach reduces to the traditional APCA method. One of the differences between the one-factor case in (5) and the multi-factor case in (10) is that the estimated factors and loadings in the one-factor case satisfy the conditions in (4), whereas those obtained in the multi-factor case may not. It is important to note that this discrepancy is only a finite sample outcome. Asymptotically, we can demonstrate that the multi-factor case still satisfies the identification conditions in (4). Similar outcomes are also observed in the sparse PCA literature, including Zou et al. (2006) and Witten et al. (2009), as well as in the sparse factor modeling frameworks discussed in Kristensen (2017).
For each sparse vector $\widehat{\mathbf{v}}_{i}$ obtained in Algorithm 2, the estimated factor process is obtained as $\underline{\widehat{\mathbf{f}}}_{i}=\sqrt{T}\widehat{\mathbf{v}}_{i}$. It is not hard to see that the number of iterations is $r$, and the $i$th iteration outputs a sparse estimator (multiplied by $\sqrt{T}$) of the $i$th column of $\mathbf{F}$}. For the optimization problem in Step 4 of Algorithm 2, by the argument in Lemma 1 of the Supplement, the matrix ${\mathbf{B}}_{i}$ in each step is a symmetric one. Therefore, we modify the truncated power method in Algorithm 1 and propose the following algorithm to obtain a sparse eigenvector for Step 4 of Algorithm 2.
@@
1:For each ${\mathbf{B}}={\mathbf{B}}_{i}$, perform a singular-value-decomposition (SVD) ${\mathbf{B}}={\mathbf{U}}{\mathbf{D}}{\mathbf{U}}^{\prime}$ and hence ${\mathbf{B}}^{1/2}={\mathbf{U}}{\mathbf{D}}^{1/2}{\mathbf{U}}^{\prime}$. Let ${\mathbf{A}}={\mathbf{B}}^{-1/2}{\mathbf{S}}{\mathbf{B}}^{-1/2}$;
@@
4:  Compute $\widetilde{\mathbf{x}}_{j}=\frac{{\mathbf{B}}^{-1/2}{\mathbf{A}}{\mathbf{x}}_{j-1}}{\|{\mathbf{A}}{\mathbf{x}}_{j-1}\|_{2}}$;
5:  Let $l_{j}=\text{supp}(\widetilde{\mathbf{x}}_{j},s)$ be the indices of $\widehat{\mathbf{x}}_{t}$ with the largest $s$ absolute values;
6:  Compute ${\mathbf{x}}_{j}^{*}=\text{Truncate}(\widetilde{\mathbf{x}}_{j},l_{j})$;
7:  Let ${\mathbf{x}}_{j}=\frac{{\mathbf{B}}^{1/2}{\mathbf{x}}_{j}^{*}}{\|{\mathbf{B}}^{1/2}{\mathbf{x}}_{j}^{*}\|_{2}}$;
8:  $j\leftarrow j+1$;
9:until ${\mathbf{x}}_{j}$ is convergent;